# Attention Is All You Need for LLM-based Code Vulnerability Localization

### v1

@@ 5.1 Experiment Setup
Big-Vul CVEFixes-J-M CVEFixes-P-M SmartFix C S-J S-P
@@
Avg Vuln line 4 3 4 1 5 7 6
TABLE I: All datasets
Baseline.
We primarily compare our approach with the direct outputs of LLMs. Following previous works [26, 42], we provide the vulnerable code to an instruction-based LLM, which outputs all lines containing vulnerabilities and repeats the process $k$ times. After $k$ iterations, the suspicious score for each line is computed. The suspicious score for line $m$ is calculated using the following formula:
@@
where $r_{i}$ represents all the vulnerable lines output by the LLM in the $i$-th iteration. In the case of ties, the lines are ranked according to the order in which they were output. For example, We set $k=10$, meaning each piece of vulnerable code is processed by the LLM to output ten times.
@@
Metrics: We use the following metrics: Top-N measures how many relevant items (e.g., correct localiaztion) appear within the top N results returned by our system. Following previous works, we use $N=1,3,5$ to evaluate our approach [42]. Precision measures the proportion of correctly identified positive instances out of all instances that were predicted as positive. Recall, also known as sensitivity or true positive rate, measures the proportion of correctly identified positive instances out of all actual positive instances. F1-Score, which is the harmonic mean of Precision and Recall, provide a balanced metric that considers both false positives and false negatives. The F1-Score provides a single metric that balances the trade-off between precision and recall. The detailed definitions and formulations are omitted here, as they are extensively covered in other works.
@@
Model: All experiments, except for RQ3 (Robustness), were conducted on the Llama-3.1-8B-Instruct [18] model, which is one of the state-of-the-art LLMs with fewer than 10b parameters.
@@
Environment: The experiments were conducted on a high-performance computing system running Ubuntu 22.04.4 LTS. The system is equipped with an NVIDIA A800 80GB PCIe GPU with CUDA version 12.4 and an AMD EPYC 7763 processor.
5.2 Experiment Results
@@
Effectiveness (RQ1): The goal of the experiment is to evaluate how accurately LOVA localizes vulnerabilities compared to existing techniques. The evaluation is conducted using four different decoding-based settings of a LLM: (i) vanilla output, (ii) few-shot chain-of-thought (CoT) output [1], (iii) Mixture-of-Agents (MoA) [44] output and (iv) rStar [45] output. In the vanilla output setting, the model directly generates the final result. In the few-shot CoT output, the model employs the few-shot CoT technique, where first reasoning examples are provided to the LLM, and then it first outputs the reasoning process before providing the final answer. This method enhances the model’s reasoning capabilities. The MoA output requires the LLM to generate multiple results and then evaluate each output by themselves. Based on these evaluations, the model will provide the final output. Using MoA can help reduce the bias introduced by the randomness of the LLM. rStar combines the reasoning and validation processes by first having the LLM generate multiple human-like reasoning actions to form different reasoning trajectories. Next, the LLM evaluates these trajectories to obtain the most mutually consistent trajectory as the final result. rStar is currently one of the state-of-the-art methods for enhancing the reasoning capabilities of LLMs based on reasoning chains construction. For each dataset, we performed 5-fold cross-validation to obtain results across the entire dataset.
@@
TABLE II: Comparison of Effectiveness Across Methods on C Language Datasets: Big-Vul and CVEFixes-C. SmartFix is excluded in this RQ as it is specific to Solidity language.
As shown in Table II, LOVA achieved the highest F1-score on both the Big-Vul and CVEFixes-C datasets compared to the baseline. Specifically, on Big-Vul, it improved by 1.5$\times$ and 2.2$\times$ over vanilla output and CoT output, respectively. On CVEFixes-C, it achieved gains of 3.4$\times$ and 5.3$\times$ over vanilla output and CoT output, respectively. While vanilla output attained the highest recall, it had the lowest precision, indicating a higher likelihood of producing false positives. By examining the model’s output, it can be seen that vanilla output often marks a large number of suspicious lines, leading to more false positives. In contrast, CoT output, after explaining the vulnerability, more precisely identifies the vulnerable lines, resulting in higher precision but lower recall. LOVA achieved the highest precision, striking a balance between precision and recall compared to the baselines, which led to the highest F1-score.
In addition to F1-score, Top-N accuracy is also a useful metric for evaluating vulnerability localization. We compare the Top-N accuracy of LOVA with both vanilla output and CoT output, and investigate how combining multiple rounds of LLM outputs affects the accuracy. The experimental results are shown in Figure 6. The x-axis represents the accuracy after combining the results of k outputs, and the y-axis represents the Top-N accuracy, where N = 1, 3, 5. LOVA demonstrates a marked improvement over both the vanilla and CoT outputs. In the Top-1, Top-3, and Top-5 categories, the vanilla output fails to generate accurate vulnerability localization results consistently. To further evaluate LOVA ’s performance in managing long contexts, we compared its effectiveness against vanilla and CoT outputs across different ranges of lines of code (LoC) using the Big-Vul and CVEFixes-C datasets, as shown in Figure 7. The results indicate a substantial decline in the accuracy of both vanilla and CoT outputs as the LoC increases, with average Top-1 accuracy dropping below 10% when the LoC exceeds 80 lines. In contrast, LOVA consistently maintains high accuracy across varying LoC ranges.
@@
Figure 6: Comparison of Top-N between LOVA and baselines
Figure 7: Top-1 Accuracy on different line ranges
@@
TABLE III: Comparison of F1 and Top-1 Scores across Models for Java, Python, and Solidity
Scalability (RQ2): One key advantage of LLMs over traditional smaller machine learning models lies in their ability to generalize across multiple languages. To evaluate the scalability, we conduct vulnerability localization in three additional languages—Python, Java, and Solidity—beyond C/C++. The results are shown in Table III. It is clear that LOVA achieved the highest Top-1 accuracy and F1-score across all datasets. On the method-level-context datasets CVEFixes-J-M and CVEFixes-P-M, LOVA improved the F1-score by up to 1.6$\times$ and 1.7$\times$ compared to the baselines, respectively. In longer contexts, LOVA achieved an impressive 14.6$\times$ improvement on the contract-level context dataset SmartFix compared to the vanilla output. The CoT output also achieved a high F1 score of 35.1 on this dataset. This may be due to the SmartFix dataset containing only a few simple types of vulnerabilities, making it easier for CoT to infer vulnerabilities. For the file-level-context datasets CVEFixes-J and CVEFixes-P, it achieved increases of up to 6.7$\times$ and 3.9$\times$, respectively, demonstrating LOVA’s remarkable ability to handle long contexts effectively.
@@
Robustness (RQ3): The robustness of LOVA across various LLM is a critical factor in assessing its general applicability. To evaluate its effectiveness, we applied LOVA to multiple state-of-the-art LLMs, including Llama-3.1-8B-Instruct [18], Mistral-7B-Instruct-v0.2 [46], and Phi-3.5-mini-instruct [47], and measured its performance across vulnerability localization tasks. The results, summarized in Table IV, show that LOVA consistently outperforms baseline methods regardless of the underlying LLM. Specifically, when applied to Mistral, LOVA improved the Top-1 accuracy by 1.1$\times$ compared to CoT outputs and achieved an F1 score that was 1.5$\times$ higher than vanilla outputs on method-level vulnerability localization tasks. On Phi, LOVA exhibited relatively low accuracy across file-level vulnerability datasets. This may be due to Phi having only 3.8b parameters, the fewest among the three models. However, LOVA still achieved a 6$\times$ increase in Top-1 accuracy compared to vanilla outputs, demonstrating its adaptability across models with varying parameter sizes. Notably, LOVA ’s performance on Llama was particularly impressive in handling longer contexts, where it outperformed CoT outputs by up to 4.9$\times$ in terms of Top-1 accuracy and showed a 3.4$\times$ increase in F1 score compared to the vanilla output on the file-level datasets. This demonstrates LOVA ’s robust capability to adapt to different LLM architectures and maintain high accuracy in diverse vulnerability localization scenarios.
@@
TABLE IV: Comparison of Methods Across C Datasets: Big-Vul and CVEFixes. SmartFix is excluded in this RQ as it is specific to Solidity language.
@@
[backgroundcolor=red!2]
Answer (RQ3). LOVA can be applied across various LLMs, as the self-attention mechanisms in different models consistently demonstrate semantic connections with the output.
Abalation Study (RQ4): To better assess the contribution of each design element, we conducted an ablation study focusing on three key components of the method: Code Line Highlight, Attention Calculation, and Vulnerability Line Localization. From this, we derived three distinct variants of LOVA: LOVA-C, LOVA-A, and LOVA-V. In LOVA-C, the code line highlighting mechanism was altered by introducing marks at specific positions that require emphasis. LOVA-A utilizes a average pooling technique for dimensionality reduction and feature extraction, compressing the attention outputs into vectors of uniform length. Lastly, LOVA-V replaces the Bi-LSTM in vulnerability localization with a multilayer perceptron (MLP) network, omitting the use of context during the process.
@@
TABLE V: Accuracy comparison across LOVA variants and C datasets
The results of the ablation study are summarized in Table V. As expected, each variant of LOVA demonstrated varying degrees of effectiveness compared to the full model. LOVA-C, which modifies the code line highlighting approach, showed a moderate reduction in performance, with the Top-1 accuracy dropping by 5% compared to the full LOVA, suggesting that the original highlighting strategy plays a critical role in precise vulnerability localization. In LOVA-V, which replaces contextual information with an MLP classifier, the accuracy dropped by 4%, indicating that contextual information plays a certain role in aiding vulnerability classification. However, LOVA-A, where attention outputs were compressed through average pooling, exhibited the most significant performance decline, with an F1 score reduced by 9%. This underscores the importance of effective dimensionality reduction and feature selection techniques for accurate vulnerability localization, as not all parts of the attention output can capture the semantic meaning relevant to identifying vulnerabilities.
@@ 6 Related Work
Vulnerability Localization. Automated vulnerability localization can significantly reduce developers’ efforts in identifying and diagnosing vulnerabilities. Traditional vulnerability localization tools, such as [39, 40], rely on predefined detectors for different vulnerabilities, limiting scalability. Deep learning-based methods[31, 30, 29, 28, 27, 25, 23, 24] focus on learning different program features to determine whether a specific line contains a vulnerability. Program features include Control Flow Graph (CFG), Program Dependency Graph (PDG), Abstract Syntax Tree (AST) and more. However, deep learning methods suffer from poor generalization. On the one hand, these methods are often limited to a specific programming language, and on the other hand, collecting various program features requires the program to be compilable or executable in order to gather relevant information, imposing additional requirements on the program structure.
@@
With the rise of LLMs, LLM-based vulnerability localization has emerged [6]. The key advantage of LLM-based approaches is their generalizability, allowing them to be applied across different types and domains of vulnerabilities. Current LLM-based methods can be broadly categorized into two types: directly output methods [2, 26] and fine-tuning-based methods [32, 26]. Our approach innovatively leverages self-attention mechanisms as indicators for vulnerability localization, significantly improving accuracy compared to directly output by LLMs, without the high cost of fine-tuning. Additionally, LOVA mitigates the performance degradation of LLM in long-context scenarios, making it better suited to real-world vulnerability localization needs.
@@
LLM Attention for Code Analysis. Self-attention mechanisms are widely used as interpretability tools for code. For code syntax interpretability, Wan et al. [43] used attention to assess whether pre-trained language models can capture syntactical relationships in source code. They observed that tokens sharing a common parent node in the Abstract Syntax Tree (AST) receive higher attention, suggesting that attention can help explain the model’s ability to capture syntactic structures. In terms of code semantic interpretability, Ma et al. [35] explored whether tokens within the same Control Dependency Graph (CDG), Control Flow Graph (CFG), or Data Dependency Graph (DDG) receive significantly increased attention. Their results indicate that many attention heads are indeed capable of capturing semantic information. Additionally, self-attention mechanisms can represent the importance of each token in a prompt relative to the output. Kou [36] examined whether the content that LLMs focus on during coding tasks aligns with human. They used the last token in the prompt and the last token in the output as query tokens, calculating the attention to each token in the prompt as a measure of the LLM’s focus during generation. We are the first to apply attention as an indicator for vulnerability localization.
7 Conclusion

### v2

@@ 5.1. Experiment Setup
Big-Vul CVEFixes-J-M CVEFixes-P-M SmartFix C J P
@@
Avg Vuln lines 4 3 4 1 5 7 6
Baseline: We primarily compare our approach with the direct outputs of LLMs. Following previous works (Zhang et al., 2024b; Kang et al., 2024b), we provide the vulnerable code to an instruction-based LLM, which outputs all lines containing vulnerabilities and repeats the process $k$ times. After $k$ iterations, the suspicious score for each line is computed. The suspicious score for line $m$ is calculated using the following formula:
@@
where $r_{i}$ represents all the vulnerable lines output by the LLM in the $i$-th iteration. In the case of ties, the lines are ranked according to the order in which they were output. We set $k=10$, meaning each piece of vulnerable code is processed by the LLM to output ten times.
@@
Metrics: We use the following metrics: Top-N measures how many relevant items (e.g., correct localization) appear within the top N results returned by our system. Following previous works, we use $N=1,3,5$ to evaluate our approach (Kang et al., 2024b). Precision measures the proportion of correctly identified positive instances out of all instances that were predicted as positive. Recall, also known as sensitivity or true positive rate, measures the proportion of correctly identified positive instances out of all actual positive instances. F1-Score, which is the harmonic mean of Precision and Recall, provides a balanced metric that considers both false positives and false negatives. The F1-Score provides a single metric that balances the trade-off between precision and recall. The detailed definitions and formulations are omitted here, as they are extensively covered in other works.
@@
Model: All experiments, except for RQ3 (Robustness), were conducted on the Llama-3.1-8B-Instruct (Touvron et al., 2023) model, which is one of the state-of-the-art LLMs with fewer than 10b parameters.
@@
Environment: The experiments were conducted on a computing system running Ubuntu 22.04.4 LTS. The system is equipped with an NVIDIA A800 80GB PCIe GPU with CUDA version 12.4 and an AMD EPYC 7763 processor.
5.2. Experiment Results
@@
Effectiveness (RQ1): The goal of the experiment is to evaluate how accurately LOVA localizes vulnerabilities compared to existing techniques. The evaluation is conducted using four different decoding-based settings of an LLM: (i) vanilla output, (ii) few-shot chain-of-thought (CoT) output (Wei et al., 2022), (iii) Mixture-of-Agents (MoA) (Wang et al., 2024) output and (iv) rStar (Qi et al., 2024) output. In the vanilla output setting, the model directly generates the final result. In the few-shot CoT output, the model utilizes the few-shot CoT technique by first presenting reasoning examples to the LLM. The model then generates a step-by-step reasoning process before providing the final answer. This method enhances the model’s reasoning capabilities. The MoA output requires the LLM to generate multiple results and then evaluate each output by themselves. Based on these evaluations, the model will provide the final output. Using MoA can help reduce the bias introduced by the randomness of the LLM. rStar combines the reasoning and validation processes by first having the LLM generate multiple human-like reasoning actions to form different reasoning trajectories. Next, the LLM evaluates these trajectories to obtain the most mutually consistent trajectory as the final result. rStar is currently one of the state-of-the-art methods for enhancing the reasoning capabilities of LLMs based on reasoning chain construction. For each dataset, we performed 5-fold cross-validation to obtain results across the entire dataset.
Table 2. Comparison of Effectiveness Across Methods on C Language Datasets: Big-Vul and CVEFixes-C. SmartFix is excluded in this RQ as it is specific to Solidity language.
@@
As shown in Table 2, LOVA achieved the highest F1-score on both the Big-Vul and CVEFixes-C datasets compared to the baseline. Specifically, on Big-Vul, it improved by 1.5$\times$ and 2.2$\times$ over vanilla output and CoT output, respectively. On CVEFixes-C, it achieved gains of 3.4$\times$ and 5.3$\times$ over vanilla output and CoT output, respectively. While vanilla output attained the highest recall, it had the lowest precision, indicating a higher likelihood of producing false positives. By examining the model’s output, it can be seen that vanilla output often marks a large number of suspicious lines, leading to more false positives. In contrast, CoT output, after explaining the vulnerability, more precisely identifies the vulnerable lines, resulting in higher precision but lower recall. Although more advanced reasoning techniques like MoA and rStar show improvement over CoT and Vanilla, they still perform significantly below LOVA. LOVA achieved the highest precision, striking a balance between precision and recall compared to the baselines, which led to the highest F1-score.
In addition to F1-score, Top-N accuracy is also a useful metric for evaluating vulnerability localization. We compare the Top-N accuracy of LOVA with both vanilla output and CoT output and investigate how combining multiple rounds of LLM outputs affects the accuracy. The experimental results are shown in Figure 6. The x-axis represents the accuracy after combining the results of k outputs, and the y-axis represents the Top-N accuracy, where N = 1, 3, 5. LOVA demonstrates a marked improvement over both the vanilla and CoT outputs. In the Top-1, Top-3, and Top-5 categories, the vanilla output fails to generate accurate vulnerability localization results consistently. To further evaluate LOVA ’s performance in managing long contexts, we compared its effectiveness against vanilla and CoT outputs across different ranges of lines of code (LoC) using the Big-Vul and CVEFixes-C datasets, as shown in Figure 7. The results indicate a substantial decline in the accuracy of both vanilla and CoT outputs as the LoC increases, with average Top-1 accuracy dropping below 10% when the LoC exceeds 80 lines. In contrast, LOVA consistently maintains high accuracy across varying LoC ranges.
@@
Figure 6. Comparison of Top-N between LOVA and baselines, where k refers to the iterations of the baseline.
Figure 7. Top-1 Accuracy on different line ranges
@@
Table 3. Comparison of F1 and Top-1 Scores Among Java, Python, and Solidity Languages
@@
Scalability (RQ2): One key advantage of LLMs over traditional smaller machine learning models lies in their ability to generalize across multiple languages. To evaluate the scalability, we conduct vulnerability localization in three additional languages—Python, Java, and Solidity—beyond C/C++. The results are shown in Table 3. It is clear that LOVA achieved the highest Top-1 accuracy and F1-score across all datasets. On the method-level-context datasets CVEFixes-J-M and CVEFixes-P-M, LOVA improved the F1-score by up to 1.6$\times$ and 1.7$\times$ compared to the baselines, respectively. In longer contexts, LOVA achieved an impressive 14.6$\times$ improvement on the contract-level context dataset SmartFix compared to the vanilla output. The CoT output also achieved a high F1 score of 35.1 on this dataset. This may be due to the SmartFix dataset containing only a few simple types of vulnerabilities, making it easier for CoT to infer vulnerabilities. For the file-level-context datasets CVEFixes-J and CVEFixes-P, it achieved increases of up to 6.7$\times$ and 3.9$\times$, respectively, demonstrating LOVA’s remarkable ability to handle long contexts effectively.
@@
Robustness (RQ3): The robustness of LOVA across various LLM is a critical factor in assessing its general applicability. To evaluate its effectiveness, we applied LOVA to multiple state-of-the-art LLMs, including Llama-3.1-8B-Instruct (Touvron et al., 2023), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023b), and Phi-3.5-mini-instruct (Abdin et al., 2024), and measured its performance across vulnerability localization tasks. The results, summarized in Table 4, show that LOVA consistently outperforms baseline methods regardless of the underlying LLM. Specifically, when applied to Mistral, LOVA improved the Top-1 accuracy by 1.1$\times$ compared to CoT outputs and achieved an F1 score that was 1.5$\times$ higher than vanilla outputs on method-level vulnerability localization tasks. On Phi, LOVA exhibited relatively low accuracy across file-level vulnerability datasets. This may be due to Phi having only 3.8b parameters, the fewest among the three models. However, LOVA still achieved a 3.4$\times$ increase in F1 score compared to CoT outputs, demonstrating its adaptability across models with varying parameter sizes. Notably, LOVA ’s performance on Llama was particularly impressive in handling longer contexts, where it outperformed CoT outputs by up to 4.9$\times$ in terms of Top-1 accuracy and showed a 3.4$\times$ increase in F1 score compared to the vanilla output on the file-level datasets. This demonstrates LOVA ’s robust capability to adapt to different LLM architectures and maintain high accuracy in diverse vulnerability localization scenarios.
Table 4. Comparison of Methods Across C Datasets: Big-Vul and CVEFixes-C. SmartFix is excluded in this RQ as it is specific to Solidity language.
@@
[backgroundcolor=red!2] Answer (RQ3). LOVA can be applied across various LLMs, as the self-attention mechanisms in different models consistently demonstrate semantic connections with the output. Ablation Study (RQ4): To better assess the contribution of each design element, we conducted an ablation study focusing on three key components of the method: Code Line Highlight, Attention Calculation, and Vulnerability Line Localization. From this, we derived three distinct variants of LOVA: LOVA-C, LOVA-A, and LOVA-V. In LOVA-C, the code line highlighting mechanism was altered by introducing marks at specific positions that require emphasis. LOVA-A utilizes an average pooling technique for dimensionality reduction and feature extraction, compressing the attention outputs into vectors of uniform length. Lastly, LOVA-V replaces the Bi-LSTM in vulnerability localization with a multilayer perception (MLP) network, omitting the use of context during the process.
Table 5. Accuracy comparison across LOVA variants.
@@
The results of the ablation study are summarized in Table 5. As expected, each variant of LOVA demonstrated varying degrees of effectiveness compared to the full model. LOVA-C, which modifies the code line highlighting approach, showed a moderate reduction in performance, with the Top-1 accuracy dropping by 5% compared to the full LOVA, suggesting that the original highlighting strategy plays a critical role in precise vulnerability localization. In LOVA-V, which replaces contextual information with an MLP classifier, the accuracy dropped by 4%, indicating that contextual information plays a certain role in aiding vulnerability classification. However, LOVA-A, where attention outputs were compressed through average pooling, exhibited the most significant performance decline, with an F1 score reduced by 9%. This underscores the importance of effective dimensionality reduction and feature selection techniques for accurate vulnerability localization, as not all parts of the attention output can capture the semantic meaning relevant to identifying vulnerabilities.
@@ 6. Related Work
Vulnerability Localization. Automated vulnerability localization can significantly reduce developers’ efforts in identifying and diagnosing vulnerabilities. Traditional vulnerability localization tools, such as (Abreu et al., 2007; Moon et al., 2014), rely on predefined detectors for different vulnerabilities, limiting scalability. Deep learning-based methods (Li et al., 2021a, b; Zhang et al., 2023; Fu et al., 2022; Fu and Tantithamthavorn, 2022; Rajput et al., 2023; Mirsky et al., 2023; Visalli et al., 2019) focus on learning different program features to determine whether a specific line contains a vulnerability. Program features include syntactic features (e.g., Abstract Syntax Tree (Li et al., 2021a, b)), semantic features (e.g., Program Dependency Graph (Li et al., 2021a; Mirsky et al., 2023)), and other program representations. However, deep learning methods suffer from poor generalization. On the one hand, these methods are often limited to a specific programming language, and on the other hand, collecting various program features requires the program to be compilable or executable to gather relevant information, imposing additional requirements on the program structure.
@@
With the rise of LLMs, LLM-based vulnerability localization has emerged (Yao et al., 2024). The key advantage of LLM-based approaches is their generalizability, allowing them to be applied across different types and domains of vulnerabilities. Current LLM-based methods can be broadly categorized into two types: direct output methods (Nong et al., 2024; Zhang et al., 2024b) and fine-tuning-based methods (Yang et al., 2024; Zhang et al., 2024b). Our approach innovatively leverages self-attention mechanisms as indicators for vulnerability localization, significantly improving accuracy compared to direct output by LLMs, without the high cost of fine-tuning. Additionally, LOVA mitigates the performance degradation of LLM in long-context scenarios, making it better suited to real-world vulnerability localization needs.
@@
LLM Attention for Code Analysis. Self-attention mechanisms are widely used as interpretability tools for code. For code syntax interpretability, Wan et al. (Wan et al., 2022) used attention to assess whether pre-trained language models can capture syntactical relationships in source code. They observed that tokens sharing a common parent node in the Abstract Syntax Tree (AST) receive higher attention, suggesting that attention can help explain the model’s ability to capture syntactic structures. In terms of code semantic interpretability, Ma et al. (Ma et al., 2024b) explored whether tokens within the same Control Dependency Graph (CDG), Control Flow Graph (CFG), or Data Dependency Graph (DDG) receive significantly increased attention. Their results indicate that many attention heads are indeed capable of capturing semantic information. Additionally, self-attention mechanisms can represent the importance of each token in a prompt relative to the output. Kou (Kou et al., 2024) examined whether the content that LLMs focus on during coding tasks aligns with humans. They used the last token in the prompt and the last token in the output as query tokens, calculating the attention to each token in the prompt as a measure of the LLM’s focus during generation. We are the first to apply attention as an indicator for vulnerability localization.
7. Conclusion